OpenStudy (anonymous):

1. Solve the absolute value equation. Check your solutions. | 5x + 13| = –7

OpenStudy (anonymous):

make two equations 1 without the absolute value signs and another with multiplying the solution by -1 5x+13=-7 5x+13=7 get the x alone 5x=-20 5x=-6 divide by five x=-4 or x=-6/5

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yep